OpenStudy (anonymous):

-5x > 45 solve the inequality?

OpenStudy (anonymous):

Divide 45 by -5, x > that quotient

OpenStudy (anonymous):

Be sure you flip the inequality symbol when dividing by a negative number. -5x>45 /-5 /-5 x<-9

OpenStudy (tkhunny):

Or, if you are like many people, and are annoyed by the flipping thing related to negative numbers, there is another way. -5x > 45 Subtract 45 -5x - 45 > 0 Add 5x -45 > 5x Divide by POSITIVE 5 -9 > x There. Same answer. Of course, it's unique, so it had better be the same answer. And no inequality flipping.

OpenStudy (anonymous):

Hey i never knew that way , thanks . Although its not that hard to flip an inequality symbol :S

OpenStudy (tkhunny):

It isn't that hard, I agree, however: 1) Far too many errors result from forgetting to do that. If your process contains a rather illogical step, you ARE likely to overlook it. 2) Most folks just memorize it and have no idea why it works.

OpenStudy (anonymous):

It has to be x > -9 else its a false statement so be careful when flipping the the sign

OpenStudy (anonymous):

tsk tsk you didnt flip the sign! & yes tkhunny your right about that one , never thought about it like that.

OpenStudy (tkhunny):

@cl10Greg No, it is the other way around. -5x > 45 Try x = -10 -5(-10) = 50 > 45 Try x = -8 -5(-8) = 40 < 45 -10 works, as -10 < -9 -8 doesn't work, as -8 > -9